OpenStudy (anonymous):

A flowerpot falls from the ledge of an apartment building. A person in an apartment below, coincidentally in possesion of a high-speed, high-precision timing system, notices that it takes 0.16 s to fall past his 4.8 m tall window. How far above the top of the window is the ledge from which the pot fell? answer needs to be in meters

OpenStudy (anonymous):

\[x= V*T+.5a *T ^{2} \] Since it fell the original velocity would be zero and your left with \[x=1/2*a*T ^{2}\] a= acceleration and the acceleration due to gravity is 9.8 and t is how long it takes to fall down so \[x=4.9*.16^{2}\] x=.12544 m so the ledge is 4.8+.125544 m tall

OpenStudy (anonymous):

It say's that that's incorrect. It also says this... In the absence of air resistance, the acceleration of the flowerpot is constant. You will need to use two constant-acceleration expressions for the distance from the ledge to both the top and bottom of the window, and solve the resulting two equations simultaneously.

OpenStudy (anonymous):

Try 0.751m as the answer if it works i'll explain it if not i'll just try again

OpenStudy (anonymous):

doesn't work

OpenStudy (anonymous):

if it helps you any they give me practice questions and tell me the answers to those. For example : A flowerpot falls from the ledge of an apartment building. A person in an apartment below, coincidentally in possesion of a high-speed, high-precision timing system, notices that it takes 0.21 s to fall past his 3.3 m tall window. How far above the top of the window is the ledge from which the pot fell? the answer is 11m

OpenStudy (anonymous):

okay i think i figured it out try 43.5 as your answer

OpenStudy (anonymous):

How'd you get that? (sorry if that sounds rude, it's just that every time I put in a wrong answer my score decreases)

OpenStudy (anonymous):

okay so the first two times i didn't read the question properly i thought it took that long to hit the ground but what i was supposed to do was \[V2=V1+aT\] a=9.8 T=0.16 \[V2^{2}=V1^{2}+2a DeltaX\] \[\sqrt{V1^{2}+2a DeltaX}=V1+aT\] simplifies to \[V1=(2DeltaX+aT ^{2})/2T\]=29.2 \[V1=aT1\] simplifies to \[T1=V1/a\] \[X=1/2*a*T1^{2}\]=43.5 use the same steps with the numbers that they give in the example problem and you should get 11

OpenStudy (anonymous):

\[DeltaX=4.8\]

OpenStudy (anonymous):

that's so confusing but fortunately it's correct. Thank you soo much for your time.

OpenStudy (anonymous):

do you think you could help me with a different question?

OpenStudy (anonymous):

yeah sure ill try to read it properly this time

OpenStudy (anonymous):

You are riding in a boat whose speed relative to the water is 5.0 m/s. The boat points at an angle of 25.8° upstream on a river flowing at 14.8 m/s. Find the time it takes for the boat to reach the opposite shore if the river is 26.5 m wide.